# Mobile-VTON: High-Fidelity On-Device Virtual Try-On

### 01 Mar 2026 v1

@@ Abstract
Virtual try-on (VTON) has recently achieved impressive visual fidelity, but most existing systems require uploading personal photos to cloud-based GPUs, raising privacy concerns and limiting on-device deployment. To address this, we present Mobile-VTON, a high-quality, privacy-preserving framework that enables fully offline virtual try-on on commodity mobile devices using only a single user image and a garment image. Mobile-VTON introduces a modular TeacherNet–GarmentNet–TryonNet (TGT) architecture that integrates knowledge distillation, garment-conditioned generation, and garment alignment into a unified pipeline optimized for on-device efficiency. Within this framework, we propose a Feature-Guided Adversarial (FGA) Distillation strategy that combines teacher supervision with adversarial learning to better match real-world image distributions. GarmentNet is trained with a trajectory-consistency loss to preserve garment semantics across diffusion steps, while TryonNet uses latent concatenation and lightweight cross-modal conditioning to enable robust garment-to-person alignment without large-scale pretraining. By combining these components, Mobile-VTON achieves high-fidelity generation with low computational overhead. Experiments on VITON-HD and DressCode at $1024{\times}768$ show that it matches or outperforms strong server-based baselines while running entirely offline. These results demonstrate that high-quality VTON is not only feasible but also practical on-device, offering a secure solution for real-world applications.
@@ 4.1 Experiment Setup
Implementation Details All experiments were conducted using 8 NVIDIA A100 GPUs (80GB). As the official implementation of SnapGen [2] was not publicly available, we re-implemented the architecture based on the descriptions provided in the original paper. To validate our reproduction, we first trained the model on ImageNet [7] at $256\times 256$ resolution for one day. The results confirmed that the reimplementation produced plausible and consistent outputs. We then adapted the model to the downstream Mobile-VTON task using a two-stage training process. In the first stage, the model was trained for 140 epochs on a merged dataset combining the upper-body subsets of DressCode [37] and VITON-HD [3]. In the second stage, it was fine-tuned for 100 additional epochs on the higher-quality DressCode subset to further improve structural consistency and synthesis quality. All training stages used a batch size of 256 and the AdamW optimizer [26] with a learning rate of $1\text{e}{-4}$ in the first stage and $5\text{e}{-5}$ in the second. The optimizer parameters were set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The overall loss function was weighted using $\lambda_{1}=1\text{e}{-2}$, $\lambda_{2}=0.5$, and $\lambda_{3}=5\text{e}{-3}$ throughout both stages.
@@ 5 Conclusion
We introduced Mobile-VTON, a unified, distillation-driven framework for high-fidelity virtual try-on directly on mobile devices. Our method leverages the proposed TGT architecture and a dual-path distillation strategy to transfer both generative behavior and perceptual priors from a large teacher diffusion model into lightweight student networks. GarmentNet extracts temporally consistent garment features, while TryonNet fuses them with person representations through structured attention and spatial latent concatenation. A lightweight Light-Adapter further enables efficient garment-text conditioning, and our training objective integrates score-based distillation, adversarial supervision, and trajectory consistency into a cohesive loss design. Mobile-VTON shows that diffusion-based try-on systems can deliver strong visual quality without sacrificing efficiency, enabling practical and personalized fashion experiences fully on-device.

### 13 Mar 2026 v4

@@ Abstract
Virtual try-on (VTON) has recently achieved impressive visual fidelity, but most existing systems require uploading personal photos to cloud-based GPUs, raising privacy concerns and limiting on-device deployment. To address this, we present Mobile-VTON, a high-quality, privacy-preserving framework that enables fully offline virtual try-on on commodity mobile devices using only a single user image and a garment image. Mobile-VTON introduces a modular TeacherNet–GarmentNet–TryonNet (TGT) architecture that integrates knowledge distillation, garment-conditioned generation, and garment alignment into a unified pipeline optimized for on-device efficiency. Within this framework, we propose a Feature-Guided Adversarial (FGA) Distillation strategy that combines teacher supervision with adversarial learning to better match real-world image distributions. GarmentNet is trained with a trajectory-consistency loss to preserve garment semantics across diffusion steps, while TryonNet uses latent concatenation and lightweight cross-modal conditioning to enable robust garment-to-person alignment without large-scale pretraining. By combining these components, Mobile-VTON achieves high-fidelity generation with low computational overhead. Experiments on VITON-HD and DressCode at $1024{\times}768$ show that it matches or outperforms strong server-based baselines while running entirely offline. These results demonstrate that high-quality VTON is not only feasible but also practical on-device, offering a secure solution for real-world applications. Code and project page are available at https://zhenchenwan.github.io/Mobile-VTON/.
@@ 4.1 Experiment Setup
Implementation Details All experiments were conducted using 8 NVIDIA A100 GPUs (80GB). As the official implementation of SnapGen [2] was not publicly available, we re-implemented the architecture based on the descriptions provided in the original paper. We adapted the model to the downstream Mobile-VTON task using a two-stage training process. In the first stage, the model was trained for 140 epochs on a merged dataset combining the upper-body subsets of DressCode [36] and VITON-HD [3]. In the second stage, it was fine-tuned for 100 additional epochs on the higher-quality DressCode subset to further improve consistency and quality. All training stages used a batch size of 256 and the AdamW optimizer [25] with a learning rate of $1\text{e}{-4}$ in the first stage and $5\text{e}{-5}$ in the second. The optimizer parameters were set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The overall loss function was weighted using $\lambda_{1}=1\text{e}{-2}$, $\lambda_{2}=0.5$, and $\lambda_{3}=5\text{e}{-3}$ throughout both stages.
@@ 5 Conclusion
We introduced Mobile-VTON, a unified distillation-driven framework that enables high-fidelity virtual try-on directly on mobile devices. Built on the proposed TGT architecture, our approach transfers both generative behaviors and perceptual priors from a large diffusion teacher to lightweight student networks. GarmentNet extracts temporally consistent garment features, while TryonNet aligns garments with person representations through structured attention and latent concatenation. A lightweight Light-Adapter further facilitates efficient garment conditioning, and the training objective integrates distillation, adversarial supervision, and trajectory consistency. Overall, Mobile-VTON demonstrates that diffusion-based try-on systems can achieve strong visual fidelity while remaining efficient enough for fully on-device deployment.